OpenStudy (anonymous):

Steve has these score on the exams for this semester: 89, 92, and 87. How much will have to make on the fourth exam to have a final average of an A (a 92% or better)?

OpenStudy (johnweldon1993):

so 89 + 92 + 87 + x ---------------- = 92 4 multiply both sides by 4 89 + 92 + 87 + x = 368 isolate x 89 + 92 + 87 + x = 368 -89 - 92 - 87 -(those totaled) x = 100 So Steve will need a 100 on the exam to get an average of 92